I have a test question that reads the question says:
SOLVE THE EQUATION
1.6x(squared)-24=0
and next time I have an answer may I put more than one question?

Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)

It is best to only have one problem per question. There are many reasons.
Do not shout! Using capital letters is shouting.
What is squared? just the 1.6 x^2 or (1.6x)^2
for now 1.6x^2-24=0
divide by 8
.2x^2-3=0
.2x^2=3
divide by .2
x^2=3/.2
x^2=30/2
x^2=15
x=+\-sqrt(15)
which will be just a little less than 4 since 4^2=16


Answer by solver91311(24713)



Looking at this it is possible that you meant that Question 1. is



In which case, divide by 6



Now proceed in one of two ways:

First method: this is the difference of two squares, so factor it according to the pattern for the difference of two squares, to wit:



In your case, recalling that



Then use the Zero Product Rule:



or



Second method:



Add 4 to both sides:



Take the square root of both sides, remembering to consider both the positive and negative roots:





Same answer either way. Next time be more careful about how you render your question so that we don't have to guess at what you mean.

John